# Adaptive Collaborative Correlation Learning-based Semi-Supervised Multi-Label Feature Selection

### v3

@@ 6.5 Ablation Study
$\quad\quad+\theta(\operatorname{Tr}(\mathbf{F}^{T}\mathbf{L}_{s}\mathbf{F})+\operatorname{Tr}(\mathbf{W}^{T}\mathbf{X}\mathbf{L}_{s}\mathbf{X}^{T}\mathbf{W})+\alpha\|\mathbf{S}\|_{F}^{2})$
@@
(3) Access-MFS-III: It performs feature selection without the adaptive collaborative learning of the instance similarity graph and the label similarity graph modules in Access-MFS.
@@
By comparing the proposed Access-MFS with its three variant methods, Table 6.4 presents the ablation experimental results in terms of AP and MaF. The experimental results regarding RL and OE can be found in Table 2 of the supplementary material. From those tables, we can see that our method outperforms methods that use only instance similarity graph learning or only label similarity graph learning (namely, Access-MFS-I and Access-MFS-II). Additionally, methods with only one type of similarity graph learning (Access-MFS-I and Access-MFS-II) perform better than those without any similarity graph learning, i.e., Access-MFS-III. This demonstrates the effectiveness of the proposed adaptive collaborative correlation learning module, which can automatically capture the local geometric structure of samples and accurately learn label correlations using existing labels. These two processes mutually enhance each other, improving feature selection performance.
@@ 7 Conclusion
To address the high dimensionality of multi-label data and the difficulty of obtaining labels for all instances, we proposed a novel semi-supervised multi-label feature selection method named Access-MFS, which uses only a subset of labeled samples. Unlike existing methods that use predefined approaches to describe either sample correlations or label correlations separately, our method enhanced the performance of semi-supervised multi-label feature selection by adaptively and collaboratively learning both sample correlations and label correlations. An generalized regression model equipped with an extended uncorrelated constraint was presented to select informative features from semi-supervised multi-label data. Meanwhile, the adaptive collaborative learning module of instance similarity graphs and label similarity graphs was integrated into the generalized regression model to preserve sample similarity in both high-dimensional and low-dimensional spaces, assign similar labels to similar samples, and establish strong label correlations for consistent predictions. An alternative iterative optimization algorithm was also developed to solve the proposed objective function. The experimental results on real-world semi-supervised multi-label data showed that Access-MFS outperformed the state-of-the-art methods.
@@ Acknowledgments
This work was supported by the Natural Science Foundation Project of Sichuan Province (No. 24NSFSC0508), the Youth Fund Project of Humanities and Social Science Research of Ministry of Education (No. 21YJCZH045), and the National Natural Science Foundation of China (No. 62176221).

### v4

@@ 6.5 Ablation Study
$\quad\quad+\theta(\operatorname{Tr}(\mathbf{W}^{T}\mathbf{X}\mathbf{L}_{s}\mathbf{X}^{T}\mathbf{W})+\operatorname{Tr}(\mathbf{F}^{T}\mathbf{L}_{s}\mathbf{F})+\alpha\|\mathbf{S}\|_{F}^{2})$
@@
(3) Access-MFS-III: It performs feature selection without the adaptive collaborative learning of the instance similarity graph and label similarity graph modules in Access-MFS.
@@
By comparing the proposed Access-MFS with its three variant methods, Table IV presents the ablation results for AP and MaF, while Table V shows the results for RL and OE. From those tables, we can see that our method outperforms methods that use only instance similarity graph learning or only label similarity graph learning (namely, Access-MFS-I and Access-MFS-II). Additionally, methods with only one type of similarity graph learning (Access-MFS-I and Access-MFS-II) perform better than those without any similarity graph learning, i.e., Access-MFS-III. This demonstrates the effectiveness of the proposed adaptive collaborative correlation learning module, which can automatically capture the local geometric structure of samples and accurately learn label correlations using existing labels. These two processes mutually enhance each other, improving feature selection performance.
TABLE IV: Means of AP and MaF for different variants of Access-MFS on eight datasets.
Metircs Methods VirusGO Medical Langlog Enron Slashdot Recreation Computer Science
AP Access-MFS $\mathbf{.8918}$ $\mathbf{.8365}$ $\mathbf{.2914}$ $\mathbf{.6198}$ $\mathbf{.5859}$ $\mathbf{.5377}$ $\mathbf{.6536}$ $\mathbf{.5027}$
Accsess-MFS-I .8624.8624 .8009.8009 .2797.2797 .5805.5805 .5453.5453 .5109.5109 .6477.6477 .4862.4862
Accsess-MFS-II .8659.8659 .7967.7967 .2871.2871 .5854.5854 .5521.5521 .5100.5100 .6453.6453 .4879.4879
Access-MFS-III .8570.8570 .7753.7753 .2652.2652 .5399.5399 .4502.4502 .4841.4841 .6349.6349 .4469.4469
MaF Access-MFS $\mathbf{.5051}$ $\mathbf{.1951}$ $\mathbf{.0092}$ $\mathbf{.0737}$ $\mathbf{.1756}$ $\mathbf{.1455}$ $\mathbf{.1152}$ $\mathbf{.0746}$
Accsess-MFS-I .4328.4328 .1770.1770 .0048.0048 .0688.0688 .1290.1290 .1199.1199 .0890.0890 .0622.0622
Accsess-MFS-II .4515.4515 .1644.1644 .0027.0027 .0682.0682 .1346.1346 .1187.1187 .0956.0956 .0605.0605
Access-MFS-III .4276.4276 .1545.1545 .0009.0009 .0610.0610 .0878.0878 .0970.0970 .0803.0803 .0340.0340
TABLE V: Means of RL and OE for different variants of Access-MFS on eight datasets.
Metircs Methods VirusGO Medical Langlog Enron Slashdot Recreation Computer Science
RL Access-MFS $\mathbf{.0711}$ $\mathbf{.0449}$ $\mathbf{.1721}$ $\mathbf{.0968}$ $\mathbf{.1337}$ $\mathbf{.1655}$ $\mathbf{.0829}$ $\mathbf{.1248}$
Accsess-MFS-I .0820.0820 .0510.0510 .1726.1726 .1076.1076 .1523.1523 .1750.1750 .0853.0853 .1294.1294
Accsess-MFS-II .0873.0873 .0500.0500 .1762.1762 .1074.1074 .1522.1522 .1743.1743 .0866.0866 .1295.1295
Access-MFS-III .0969.0969 .0563.0563 .1794.1794 .1115.1115 .1735.1735 .1793.1793 .0878.0878 .1379.1379
OE Access-MFS $\mathbf{.1677}$ $\mathbf{.2007}$ $\mathbf{.8395}$ $\mathbf{.3203}$ $\mathbf{.5354}$ $\mathbf{.5900}$ $\mathbf{.4281}$ $\mathbf{.6157}$
Accsess-MFS-I .2113.2113 .2515.2515 .8491.8491 .3317.3317 .5846.5846 .6263.6263 .4339.4339 .6413.6413
Accsess-MFS-II .1984.1984 .2590.2590 .8422.8422 .3389.3389 .5740.5740 .6285.6285 .4387.4387 .6358.6358
Access-MFS-III .2226.2226 .2891.2891 .8731.8731 .3408.3408 .7203.7203 .6661.6661 .4482.4482 .6917.6917
@@ 7 Conclusion
To address the high dimensionality of multi-label data and the difficulty of obtaining labels for all instances, we proposed a novel semi-supervised multi-label feature selection method named Access-MFS, which uses only a subset of labeled samples. Unlike existing methods that use predefined approaches to describe either sample similarity or label similarity separately, our method enhanced the performance of semi-supervised multi-label feature selection by adaptively and collaboratively learning both sample and label similarity correlations. An generalized regression model equipped with an extended uncorrelated constraint was presented to select informative features from semi-supervised multi-label data. Meanwhile, the adaptive collaborative learning module of instance similarity graphs and label similarity graphs was integrated into the generalized regression model to preserve sample similarity in both high-dimensional and low-dimensional spaces, assign similar labels to similar samples, and establish strong label correlations for consistent predictions. An alternative iterative optimization algorithm was also developed to solve the proposed objective function. The experimental results on real-world semi-supervised multi-label data showed that Access-MFS outperformed the state-of-the-art methods.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 72495122), the Natural Science Foundation Project of Sichuan Province (No. 2024NSFSC0504), and the Youth Fund Project of Humanities and Social Science Research of Ministry of Education (No. 21YJCZH045).